OpenStudy (abbycross167):

Can someone please help me answer this question? Please? I would really appreciate it... Question~ Solve the systems using the elimination method. y = 2x - 1 y = -x + 7

OpenStudy (abbycross167):

My Answer~ Y=x-1A -2x+y=-1 X+y=7 = -2x+y=-1 Multiply by 1 X+y=7 multiply by 2 = 2x-1=-x+7 3x-1=7 3x=8 X=8/3 Y=13/3 Final answer~ (8/3, 13,3)

OpenStudy (abbycross167):

What my teacher said was wrong~ -2x+y=-1 x+y=7 good! Yes, multiply by 2, but you didn't do it!?!

OpenStudy (abbycross167):

can someone please show be how to fix my answer?